OpenStudy (anonymous):

Find the domain of this function: f(x) = square root x^2 -16

OpenStudy (jdoe0001):

\(\large f(x) = \sqrt{x^2-16}\) the domain of that function, or that is, the values "x" can take safely are the ones that will NOT make the radical amount, negative

OpenStudy (anonymous):

Eh, didn't mean to ruin this for you =S deleted mine

OpenStudy (jdoe0001):

hehe... didn't have to, a helping hand is always welcome

OpenStudy (anonymous):

So should I post it or not? :S

OpenStudy (jdoe0001):

\(\bf f(x) = \sqrt{x^2-16}\\ \quad \\ x = 4\qquad \sqrt{(4)^2-16}\implies\sqrt{16-16}\implies 0\\ \quad \\ x = -4\qquad \sqrt{(-4)^2-16}\implies\sqrt{16-16}\implies 0\\\quad \\ x = \pm 3\qquad \sqrt{(\pm 3)^2-16}\implies \sqrt{9-16}\implies\color{red}{\sqrt{-7}}\\ \quad \\ x = \pm 5\qquad \sqrt{(\pm 5)^2-16}\implies \sqrt{25-16}\implies \sqrt{9}\implies \pm3\) so as you can see, 4 or -4 works well, no negatives you go below that, you get a negative value you go above that, you're ok so that's the domain

OpenStudy (jdoe0001):

@pitamar sure, post away

OpenStudy (anonymous):

\[ f(x) = \sqrt{x^2 - 16} \\ x^2 - 16 \ge 0 \\ x^2 \ge 16 \\ |x| \ge 4 \\ x \ge 4 \text{ or } x \le -4 \] Small fix also.. hehe..